OpenStudy (anonymous):

if 2^(x+3)=32 then the value of 3^(x+1)=

Parth (parthkohli):

When you have the same base, you may equate the exponent. \( \color{Black}{\Rightarrow 2^{x + 3} = 2^5 }\) \( \color{Black}{\Rightarrow x+3 = 5 }\) Solve for \(x\).

OpenStudy (jiteshmeghwal9):

Therefore,x=5-3=2.